OpenStudy (anonymous):

HELP!!!! Solve the system of equations: 2x+2y+z=-7 -x+2y+4z=-7 x+y+z=-5 The solution is ( _, _, _,)?

OpenStudy (anonymous):

use elimination always the quickest :D

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

ok but i shall use augmented matrices it is quicker if i dont have to keep writing the +,= and x y z 2x+2y+z=-7 -x+2y+4z=-7 x+y+z=-5 is